OpenStudy (jonathanvaldes98):

change the equation into slope-intercept form i need to know how to do it to finish the others so dont relly need the answer but how to do it –3x + 2y = 8

OpenStudy (ivettef365):

just isolate the y, so start by transferring the -3x to the other side, when you move from one side to another remember to change the sign do you know hot to ?

OpenStudy (jonathanvaldes98):

2y=11x

OpenStudy (amistre64):

you cant combine unlike terms ....

OpenStudy (amistre64):

-3x + 2y = 8 +3x +3x --------------- 0x + 2y = 3x + 8 2y = 3x + 8 ; and continue to solve for the y

OpenStudy (amistre64):

for those just learning the process, i do not think refering to it as "moving" to the other side presents an accurate view of the mechanics involved.

OpenStudy (jonathanvaldes98):

so am i solving it like a 2-step equation

OpenStudy (ivettef365):

maybe @amistre64 can help, he got better terms to explain

OpenStudy (amistre64):

yes, similar to a 2 step equation. All the algebraic properties are still usable, but instead of solving for a specific value of y, we are expressing it as a rule that defines multiple values of y

OpenStudy (amistre64):

the solution set, produces a line on a graph instead of a singular point.